# Uncertainty-Aware Critic Augmentation for Hierarchical Multi-Agent EV Charging Control

### v1

@@ Abstract
The advanced bidirectional EV charging and discharging technology, aimed at supporting grid stability and emergency operations, has driven a growing interest in workplace applications. It not only effectively reduces electricity expenses but also enhances the resilience of handling practical issues, such as peak power limitation, fluctuating energy prices, and unpredictable EV departures. However, existing EV charging strategies have yet to fully consider these factors in a way that benefits both office buildings and EV users simultaneously. To address these issues, we propose HUCA, a novel real-time charging control for regulating energy demands for both the building and electric vehicles. HUCA employs hierarchical actor-critic networks to dynamically reduce electricity costs in buildings, accounting for the needs of EV charging in the dynamic pricing scenario. To tackle the uncertain EV departures, a new critic augmentation is introduced to account for departure uncertainties in evaluating the charging decisions, while maintaining the robustness of the charging control. Experiments on real-world electricity datasets under both simulated certain and uncertain departure scenarios demonstrate that HUCA outperforms baselines in terms of total electricity costs while maintaining competitive performance in fulfilling EV charging requirements. A case study also manifests that HUCA effectively balances energy supply between the building and EVs based on real-time information.
@@ 1 Introduction
Background. Electric Vehicle (EV) sales have seen significant increases in recent years, largely fueled by a worldwide emphasis on sustainability. According to a report by CBRE Group (Coldwell Banker Richard Ellis Group) Incorporated [4], workplace charging sessions increased at twice the rate of new charging station installations in 2023. This highlights the growing interest for EV charging options in the workplace.
@@
Meanwhile, recent advances in bidirectional charging systems [20, 9, 25], which support both Grid-to-Vehicle (G2V) and Vehicle-to-Grid (V2G) power flows, have shown the potential to enhance flexibility with regard to charging stability and emergency situations. The effectiveness of bidirectional charging systems has inspired us to explore how better scheduling of G2V and V2G timings can balance the demand for EV charging and reduce overall electricity costs. In particular, EVs often remain stationary for long periods at office locations, as expected. This prolonged parking time allows for more flexible charging strategies, enabling energy transfer between EVs and office building. An example of a workplace charging system is shown in Figure 1. Upon arrival, each EV is coordinated by the charging system, with users initially supplying details like their charging demands and planned departure time. Based on this, the system decides whether to discharge energy from the EV to the building (using V2G) during high-price periods or to charge the EV from the grid to meet its demand before departure (using G2V). Either the V2G or the G2V option can be determined on-the-fly according to the optimal decision-making criteria.
@@
Challenges. For real-time charging control of EVs in various scenarios, previous studies [21, 6] have explored the use of multi-agent reinforcement learning (MARL) techniques to regulate EV charging and discharging actions. However, most existing approaches fail to consider real-world dynamic factors, such as dynamic energy prices and the possibility that EV users may depart earlier than the expected time, which complicate determining optimal control strategies for each EV. Moreover, to avert transformer overloads that could destabilize the power grid [26], it is necessary to impose charging power limits, thereby further restricting the management of EV charging. These dynamics and limitations make it challenging to balance the energy supply between the building and all EVs while minimizing electricity costs. It is important to note that improper management of charging can lead to notably increased electricity expenses, as power companies might levy extra charges due to overconsumption of energy [19].
@@
Proposed Method. To tackle these challenges, we propose HUCA (Hierarchical Multi-Agent Control with Uncertainty-Aware Critic Augmentation), a novel framework designed for real-time charging control. HUCA operates under dynamic pricing and penalty mechanisms, consistent with practical situations. A new hierarchical multi-agent framework is devised to balance the energy supply between buildings and EVs in dynamic and uncertain environments. The HUCA framework consists of two levels of control: a high-level agent and multiple low-level agents. The high-level agent determines whether to charge or discharge EVs, which is a collective decision for all EVs. In light of the high-level decision, the low-level agents collaboratively and dynamically modulate the individual charging (or discharging) power level for each EV, ensuring they stay within the specified charging power limits, with the goal of satisfying anticipated demands and avoiding transformer overload. The action spaces of high-level and low-level agents are respectively discrete and continuous due to different goals (detailed later).
@@
A key innovation of HUCA lies in its capability to handle EVs that deviate from their expected departure times, whether leaving earlier or later. As these departures are unpredictable, rather than letting this variability spread throughout the low-level networks and interfere with future decisions, we propose an uncertainty-aware critic augmentation method. This mechanism is designed to (i) consider uncertainties when assessing the quality of charging decisions by low-level agents and (ii) limit the direct effect of uncertainties to the present time slot. Therefore, HUCA dynamically adapt to arbitrary EV behaviors while maintaining robust charging control. With this design, HUCA is able to minimize the electricity costs of the building while striving to fulfill EV charging demands in a dynamic environment.
@@
  • Uncertainties and Limitations Handling: In HUCA, the high-level and low-level agents determine optimal actions under dynamic pricing and power stability constraints. Additionally, an uncertainty-aware augmentation is designed to adaptively adjust decisions to handle uncertain EV departures.
  • Practical Effectiveness: We evaluate HUCA through analysis of real-world electricity datasets, incorporating simulated EV behaviors in scenarios with both certain and uncertain departure times. Results demonstrate that HUCA achieves the lowest electricity costs while maintaining competitive performance in fulfilling EV charging requirements.
@@ 2 Preliminaries
$\left\{\begin{matrix}\mathcal{P}^{\max}_{i,t}=\min\bigl{(}P^{\text{pile}}_{t},(SoC^{UB}_{t}-SoC^{j}_{t-1}\cdot C^{i}\cdot\eta^{*})\bigr{)}\\ \mathcal{P}^{\min}_{i,t}=\max\bigl{(}-P^{\text{pile}}_{t},(SoC^{LB}_{t}-SoC^{j}_{t-1}\cdot C^{j}\cdot\eta^{*})\bigr{)}\end{matrix}\right.,\vskip-3.00003pt$
@@
It is generally recognized that the single-agent DDPG model is insufficient for optimizing multiple objectives, as our work aims to benefit both the office building and EVs. As such, we enhance the DDPG framework and model the situation as a hierarchical MDP to simultaneously optimize multiple objectives, aiming to find the most effective bidirectional charging strategy that reduces building overall electricity costs while considering EV charging needs.
@@ 3 The HUCA Framework
(3.2) $\hskip-2.84544ptr_{h}=\kappa\cdot\Bigl{(}\underbrace{-\sum_{t^{\prime}=1}^{t}p_{t^{\prime}}\cdot L_{t^{\prime}}-\sum_{t^{\prime}=1}^{t}\max(0,\Delta L_{t^{\prime}})\cdot\varphi}_{\text{electricity cost term}}\Bigr{)}$
@@
(3.3) $\hskip-2.84544pt\mathcal{L}(\theta^{Q}_{h})=\mathop{\mathbb{E}}_{s_{h},a_{h},r_{h},s^{\prime}_{h}\sim\mathcal{D}_{h}}\biggl{[}\Bigl{(}Q_{h}(s_{h},a_{h})-y_{h}\Bigr{)}^{2}\biggr{]},$
$\hskip-2.84544pty_{h}=r_{h}+\gamma\bar{Q}_{h}(s^{\prime}_{h},a^{\prime}_{h})|_{a^{\prime}_{h}=\mu_{h}(s^{\prime}_{h})},$
@@
(3.5) $\hskip-2.84544pt\mathcal{L}(\theta^{Q,i}_{l})=\mathop{\mathbb{E}}_{\textbf{x},\textbf{a},\textbf{r},\textbf{x}^{\prime}\sim\bigcup_{c_{i}\in\mathcal{C}}\{\mathcal{D}^{i}_{l}\}}\biggl{[}\Bigl{(}Q^{i}_{l}(\textbf{x},a^{1}_{l},.,a^{N}_{l})-y^{i}_{l}\Bigr{)}^{2}\biggr{]},$
$\hskip-2.84544pty^{i}_{l}=r^{i}_{l}+\gamma\bar{Q}^{i}_{l}(\textbf{x}^{\prime},a^{\prime 1}_{l},.,a^{\prime N}_{l})|_{a^{\prime j}_{l}=\mu^{j}_{l}(s^{\prime j}_{l})},$
@@
(3.6) $\hskip-8.5359pt\mathcal{Q}^{i}_{l}(\textbf{x},\textbf{a})=Q^{i}_{l}(\textbf{x},a^{1}_{l},.,a^{N}_{l})\cdot\biggl{(}1-$
@@
where $\sigma(\cdot)$ is the sigmoid function, and $\delta$ is set to 0.5. In this setting, the action space of $a^{i}_{l}$ is constrained to [0,0.5)[0,0.5)[ 0 , 0.5 ) while the discrete high-level action is “discharging” ($a^{\text{disc}}_{h}=0$). In contrast, $a^{i}_{l}$ is limited to [0.5,1][0.5,1][ 0.5 , 1 ] when the high-level action is “charging” ($a^{\text{disc}}_{h}=1$).
@@ 4 Experimental Results
Table 2: Performance comparison with different number of charging piles (CP). Except of the “OPT” method, the best and second best results for “Penalty Cost” ($\downarrow$) and “Total Cost” ($\downarrow$) are in bold and underlined, respectively. Cells in gray highlight the results that are greater than or equal to the median value for “SoC Maintenance” ($\uparrow$), “SoC Fulfillment” ($\uparrow$) and “User Satisfaction” ($\uparrow$).
Certain Departure Scenario
Metrics CP Num. Methods
OPT DDPG IQL VDN MADDPG HUCA
@@
Penalty Cost (USD) 20 0.00 7818.52 5643.15 2271.32 1869.38 1618.56
Total Cost (USD) 5860.05 13929.11 11798.68 8413.61 7952.69 7782.13
Uncertain Departure Scenario
Metrics CP Num. Methods
OPT DDPG IQL VDN MADDPG HUCA
@@
Total Cost (USD) 5691.96 10398.77 7479.04 6214.53 6087.48 5998.77
SoC Fulfillment (%) 53.77 54.05 46.04 40.34 27.10 41.67
SoC Maintenance (%) 42.52 -10.58 129.01 -120.61 11.87 12.80
User Satisfaction (%) 48.14 21.73 87.52 -40.13 19.48 27.23
Penalty Cost (USD) 20 0.00 5010.81 2002.12 966.51 931.01 920.43
Total Cost (USD) 5760.05 10823.07 7815.99 6721.82 6562.80 6534.15
SoC Fulfillment (%) 56.06 55.27 47.94 48.88 40.34 46.54
SoC Maintenance (%) 32.77 12.56 0.90 91.96 -1.16 10.01
User Satisfaction (%) 44.41 33.91 24.42 70.42 19.59 28.27
Table 3: Ablation study results for the uncertain departure scenario. (SoC fulfillment and maintenance scores are abbreviated as “Fulfill.” and “Main.”, respectively.)
Methods CP Num. = 10
Penalty $\downarrow$ Total $\downarrow$ Fulfill. $\uparrow$ Main. $\uparrow$
HUCA 355.84 5998.77 41.67 % 12.80%
w/o C.A. 355.87 5999.64 28.43% 12.07%
w/o H. 446.12 6087.23 40.44% 12.21%
w/o All 447.92 6087.64 27.71% 11.91%
Methods CP Num. = 20
Penalty $\downarrow$ Total $\downarrow$ Fulfill. $\uparrow$ Main. $\uparrow$
HUCA 920.43 6534.15 46.54% 10.01%
w/o C.A. 921.34 6535.14 41.34% 0.10%
w/o H. 930.51 6562.92 46.22% 9.98%
w/o All 931.73 6564.49 40.23% 0.07%
@@
As shown in Table 2, except for the ”OPT” method (best-case result), HUCA achieves the lowest penalty and total costs among methods that do not rely on future information for charging control across both departure scenarios. In addition, its total cost is comparable to OPT, manifesting its effectiveness in minimizing costs. Although the SoC fulfillment, maintenance, and user satisfaction scores of HUCA are not the best, it still achieves a comparable results to all the other baselines. Note that combining the above observations, HUCA actually strikes the best balance between the trade-off of electricity costs and fulfilling user demands by introducing the hierarchical control and the uncertainty-aware critic augmentation.
Among the multi-agent models, VDN and MADDPG come closest to HUCA in terms of cost. However, the SoC fulfillment scores of MADDPG are much lower than those of HUCA across different numbers of charging piles. While VDN achieves higher SoC fulfillment and maintenance scores than HUCA with 20 charging piles, its scores are significantly lower with 10 charging piles. The maintenance score of VDN even drops to a negative value (-120.61%), indicating that under MADDPG and VDN control, EVs’ SoCs are not maintained at arrival levels. Therefore, EVs primarily serve as power providers rather than being effectively charged, resulting in additional expenses for users and putting them at a disadvantage.
@@
To validate each component of HUCA, we conduct an ablation study by removing the uncertainty-aware critic augmentation (w/o C.A.), the high-level agent (w/o H.), and both of them (w/o All). The results are shown in Table 3. Compared to HUCA, removing the critic augmentation remains similar penalty and total costs, but decreases the SoC fulfillment score significantly. On the other hand, removing the high-level agent increases both penalty and total costs. These findings reveal that the hierarchical structure and critic augmentation of HUCA can help to achieve the lowest costs and effectively balance the energy supply between the office building and EVs.
Furthermore, to assess the impact of the uncertainty term in HUCA for handling uncertain departures, Figure 3 compares the SoC maintenance and fulfillment scores under various uncertainty coefficient $\rho$, defined in Eq. 3.6. Although the SoC maintenance score does not change significantly, the fulfillment score improves more noticeably when $\rho$ exceeds $10^{-1}$. The outcome indicates that increasing the importance of the uncertainty component improves the charging control’s ability to fulfill EV users’ expected State of Charge (SoC) requirements.

### v3

@@ Abstract
The advanced bidirectional EV charging and discharging technology, aimed at supporting grid stability and emergency operations, has driven a growing interest in workplace applications. It not only reduces electricity expenses but also enhances the resilience in handling practical matters, such as peak power limitation, fluctuating energy prices, and unpredictable EV departures. Considering these factors systematically can benefit energy efficiency in office buildings and for EV users simultaneously. To employ AI to address these issues, we propose HUCA, a novel real-time charging control for regulating energy demands for both the building and EVs. HUCA employs hierarchical actor-critic networks to dynamically reduce electricity costs in buildings, accounting for the needs of EV charging in the dynamic pricing scenario. To tackle the uncertain EV departures, we introduce a new critic augmentation to account for departure uncertainties in evaluating the charging decisions, while maintaining the robustness of the charging control. Experiments on real-world electricity datasets under both simulated certain and uncertain departure scenarios demonstrate that HUCA outperforms baselines in terms of total electricity costs while maintaining competitive performance in fulfilling EV charging requirements. A case study also manifests that HUCA effectively balances energy supply between the building and EVs based on real-time information, showcasing its potential as a key AI-driven solution for vehicle charging control.
@@ 1 Introduction
Background. The rapid growth in Electric Vehicle (EV) adoption, driven by global sustainability efforts, has led to a surge in demand for advanced charging solutions. According to a report by CBRE Group (Coldwell Banker Richard Ellis Group) Incorporated [4], workplace charging sessions increased at twice the rate of new charging station installations in 2023. This highlights the growing interest for EV charging options in the workplace. Meanwhile, recent advances in bidirectional charging systems [19] [24], which support both Grid-to-Vehicle (G2V) and Vehicle-to-Grid (V2G) power flows, have shown the potential to enhance flexibility with regard to charging stability and emergency situations. The effectiveness of bidirectional charging systems has inspired us to explore how better scheduling of G2V and V2G timings can balance the demand for EV charging and reduce overall electricity costs.
@@
In particular, EVs often remain stationary for long periods at office locations, as expected. This prolonged parking time allows for more flexible charging strategies, enabling energy transfer between EVs and office building. An example of a workplace charging system is shown in Figure 1. Upon arrival, each EV is coordinated by the charging system, with users first providing information such as their charging requirements and anticipated departure time. With such information, the system decides whether to discharge energy from the EV to the building (using V2G) during high-price periods or to charge the EV from the grid to meet its demand before departure (using G2V). Either the V2G or the G2V option can be determined on-the-fly according to the optimal decision-making criteria.
@@
Challenges. For real-time charging control of EVs in various scenarios, previous studies [20] [6] have explored the use of multi-agent reinforcement learning (MARL) techniques to regulate EV charging and discharging actions. However, most existing approaches fail to consider real-world dynamic factors, such as dynamic energy prices and the possibility that EV users may depart earlier than the expected time, which complicate determining optimal control strategies for each EV. Moreover, to avert transformer overloads that could destabilize the power grid [25], it is necessary to impose charging power limits, thereby further complicating the management of EV charging. These dynamics and limitations pose significant challenges in balancing the energy supply between the building and EVs while minimizing electricity costs. It is crucial to recognize that managing charging improperly could result in considerably higher electricity bills, as power companies will levy extra charges due to overconsumption of energy [18].
@@
Proposed Method. To tackle these challenges, we propose HUCA (Hierarchical Multi-Agent Control with Uncertainty-Aware Critic Augmentation), a novel framework designed for real-time charging control. HUCA operates under dynamic pricing and penalty mechanisms, consistent with practical situations. A new hierarchical multi-agent framework is devised to balance the energy supply between the building and EVs in dynamic and uncertain environments. The HUCA framework consists of two levels of control: a high-level agent and multiple low-level agents. The high-level agent determines whether to charge or discharge EVs, which is a collective decision for all EVs. Based on the high-level decision, the low-level agents collaboratively and dynamically modulate the individual charging (or discharging) power level for each EV, ensuring they stay within the specified charging power limits, with the goal of satisfying anticipated demands and avoiding transformer overload.
@@
A key innovation of HUCA lies in its capability to handle EVs that deviate from their expected departure times. These unpredictable departures introduce uncertainty, potentially disrupting low-level agents’ charging decisions. To address this, we propose an uncertainty-aware critic augmentation mechanism, which assesses the likelihood of departure deviation and adjusts the assessments of low-level agents’ charging decisions accordingly. This mechanism achieves two objectives: (i) accounting for uncertainties into the assessment of low-level agents’ decisions, and (ii) limiting the direct effect of uncertainties to the present time slot, ensuring the robustness of future decision assessments. Therefore, HUCA dynamically adapts to arbitrary EV behaviors while maintaining robust charging control. HUCA is designed to minimize the electricity costs of the building while striving to fulfill EV charging demands in a dynamic environment.
@@
  • Uncertainties and Limitations Handling: High-level and low-level agents determine optimal actions under dynamic pricing and power stability constraints. In addition, an uncertainty-aware augmentation is designed to adaptively adjust decisions to handle uncertain EV departures.
  • Practical Effectiveness: We evaluate HUCA through analysis of real-world datasets, incorporating simulated EV behaviors in scenarios with both certain and uncertain departure times. Results demonstrate that HUCA achieves the lowest electricity costs while maintaining competitive performance in fulfilling EV charging requirements, which shows its potential as an AI-driven solution for intelligent vehicle charging control.
@@ 2 Preliminaries
$\left\{\begin{matrix}\mathcal{P}^{\max}_{i,t}=\min\bigl{(}P^{\text{pile}}_{t},(SoC^{UB}_{t}-SoC^{j}_{t-1}\cdot C^{i}\cdot\eta^{*})\bigr{)}\\ \mathcal{P}^{\min}_{i,t}=\max\bigl{(}-P^{\text{pile}}_{t},(SoC^{LB}_{t}-SoC^{j}_{t-1}\cdot C^{j}\cdot\eta^{*})\bigr{)}\end{matrix}\right.,$
@@
Our work aims to benefit both the office building and EVs by enhancing the original single-agent DDPG framework. We model the problem as a hierarchical MDP to optimize multiple objectives simultaneously, aiming to find the most effective bidirectional charging strategy that reduces building overall electricity costs while considering EV charging needs.
@@ 3 The HUCA Framework
(3.2) $r_{h}=\kappa\cdot\Bigl{(}\underbrace{-\sum_{t^{\prime}=1}^{t}p_{t^{\prime}}\cdot L_{t^{\prime}}-\sum_{t^{\prime}=1}^{t}\max(0,\Delta L_{t^{\prime}})\cdot\varphi}_{\text{electricity cost term}}\Bigr{)}$
@@
(3.3) $\mathcal{L}(\theta^{Q}_{h})=\mathop{\mathbb{E}}_{s_{h},a_{h},r_{h},s^{\prime}_{h}\sim\mathcal{D}_{h}}\biggl{[}\Bigl{(}Q_{h}(s_{h},a_{h})-y_{h}\Bigr{)}^{2}\biggr{]},$
$y_{h}=r_{h}+\gamma\bar{Q}_{h}(s^{\prime}_{h},a^{\prime}_{h})|_{a^{\prime}_{h}=\mu_{h}(s^{\prime}_{h})},$
@@
(3.5) $\mathcal{L}(\theta^{Q,i}_{l})=\mathop{\mathbb{E}}_{\textbf{x},\textbf{a},\textbf{r},\textbf{x}^{\prime}\sim\bigcup_{c_{i}\in\mathcal{C}}\{\mathcal{D}^{i}_{l}\}}\biggl{[}\Bigl{(}Q^{i}_{l}(\textbf{x},a^{1}_{l},...,a^{N}_{l})-y^{i}_{l}\Bigr{)}^{2}\biggr{]},$
$y^{i}_{l}=r^{i}_{l}+\gamma\bar{Q}^{i}_{l}(\textbf{x}^{\prime},a^{\prime 1}_{l},...,a^{\prime N}_{l})|_{a^{\prime j}_{l}=\mu^{j}_{l}(s^{\prime j}_{l})},$
@@
(3.6) $\mathcal{Q}^{i}_{l}(\textbf{x},\textbf{a})=Q^{i}_{l}(\textbf{x},a^{1}_{l},...,a^{N}_{l})\cdot\biggl{(}1-$
@@
where $\sigma(\cdot)$ is the sigmoid function, and $\delta$ is set to 0.5. In this setting, the action space of $a^{i}_{l}$ is constrained to $[0,0.5)$ while the discrete high-level action is “discharging” ($a^{\text{disc}}_{h}=0$). In contrast, $a^{i}_{l}$ is limited to $[0.5,1]$ when the high-level action is “charging” ($a^{\text{disc}}_{h}=1$).
@@ 4 Experimental Results
Table 2: Performance comparison (HUCA performance) with different number of charging piles (CP). Except of the “OPT” method, the best and second best results for “Penalty Cost” ($\downarrow$) and “Total Cost” ($\downarrow$) are in bold and underlined, respectively. Cells in gray highlight the results that are greater than or equal to the median value for “SoC Maintenance” ($\uparrow$), “SoC Fulfillment” ($\uparrow$) and “User Satisfaction” ($\uparrow$).
Certain Departure Scenario Metric CP OPT DDPG IQL VDN MADDPG HUCA
@@
Penalty Cost (USD) 20 0.00 7818.52 5643.15 2271.32 1869.38 1643.47
Total Cost (USD) 5860.05 13929.11 11798.68 8413.61 7952.69 7728.57
Uncertain Departure Scenario Metric CP OPT DDPG IQL VDN MADDPG HUCA
@@
Total Cost (USD) 5691.96 10398.77 7479.04 6214.53 6087.48 5993.78
SoC Fulfillment (%) 53.77 54.05 46.04 40.34 27.10 40.77
SoC Maintenance (%) 42.52 -10.58 129.01 -120.61 11.87 45.40
User Satisfaction (%) 48.14 21.73 87.52 -40.13 19.48 43.08
Penalty Cost (USD) 20 0.00 5010.81 2002.12 966.51 931.01 354.90
Total Cost (USD) 5760.05 10823.07 7815.99 6721.82 6562.80 6013.52
SoC Fulfillment (%) 56.06 55.27 47.94 48.88 40.34 42.44
SoC Maintenance (%) 32.77 12.56 0.90 91.96 -1.16 124.41
User Satisfaction (%) 44.41 33.91 24.42 70.42 19.59 83.42
Table 3: Ablation study results for the uncertain departure scenario. (SoC fulfillment and maintenance scores are abbreviated as “Ful.” and “Main.”, respectively.)
CP Method Penalty $\downarrow$ Total $\downarrow$ Ful. $\uparrow$ Main. $\uparrow$
10 HUCA 355.84 5993.78 40.77 % 45.40%
w/o C.A. 2015.62 7761.55 62.49% 124.01%
w/o H. 2707.71 8483.96 63.10% 31.91%
w/o Either 723.46 6393.26 44.53% -12.53%
CP Method Penalty $\downarrow$ Total $\downarrow$ Ful. $\uparrow$ Main. $\uparrow$
20 HUCA 354.90 6013.52 42.44% 124.41%
w/o C.A. 9362.87 15309.23 69.73% 209.70%
w/o H. 3461.37 9261.72 58.61% 7.69%
w/o Either 618.46 6381.87 53.26% 43.02%
@@
As shown in Table 2, except for the ”OPT” method (best-case result), HUCA achieves the lowest penalty and total costs that do not rely on future information for charging control across both departure scenarios. In addition, its total cost is comparable to OPT, manifesting its effectiveness in minimizing costs. Although the SoC fulfillment, maintenance, and user satisfaction scores of HUCA are not the best, it still achieves a comparable results to all the other baselines. Note that combining the above observations, HUCA actually strikes the best balance between the trade-off of electricity costs and fulfilling user demands by introducing the hierarchical control and the uncertainty-aware critic augmentation.
Among the multi-agent models, VDN and MADDPG come closest to HUCA in terms of cost. However, the user satisfaction scores of MADDPG are much lower than those of HUCA across different numbers of charging piles. While VDN achieves higher SoC fulfillment score than HUCA with 20 charging piles, its maintenance scores are significantly lower. With 10 charging piles, VDN’s maintenance score even drops to a negative value (-120.61%), indicating that under MADDPG and VDN control, EVs’ SoCs are not maintained at arrival levels. Therefore, EVs primarily serve as power providers rather than being effectively charged, resulting in additional expenses for users.
@@
We conduct an ablation study by removing the uncertainty-aware critic augmentation (w/o C.A.), the high-level agent (w/o H.), and both of them (w/o Either). The results are shown in Table 3. Compared to HUCA, removing either the critic augmentation or the high-level agent increases SoC fulfillment and maintenance scores. However, this also leads to a significant rise in both penalty and total costs. On the other hand, removing both components results in a slight increase in costs but causes a significant drop in SoC maintenance scores. These findings indicate that the hierarchical structure and critic augmentation in HUCA complement each other, and their combined use is essential for achieving the lowest costs while effectively balancing energy supply between the office building and EVs.
Furthermore, to assess the impact of the uncertainty term in HUCA for handling uncertain departures, Figure 3 compares total cost and the SoC fulfillment under various uncertainty coefficient $\rho$ (Eq. 3.6). As $\rho$ increases, total cost decreases significantly. Meanwhile, setting $\rho$ to its lowest value ($10^{-2}$) improves SoC fulfillment but comes at the cost of significantly higher total costs. These findings suggest that $\rho$ should not be set too low to ensure a better trade-off between cost and user charging demand. Properly tuning $\rho$ can further improve this balance, enhancing the overall effectiveness of HUCA.